A man start from a point A and walks a distance of 3.0km due Northeast to point B. From point B he then walk 4.0 km due Southeast to point X.
Calculate the shortest distance between A and X.
1
Expert's answer
2019-10-03T09:37:12-0400

Let's choose the "Ox" axis along the east direction and "Oy" axis along the north direction and let "\varphi" be the angle with axis "Ox". Then unit vector in direction with angle "\varphi" will have coordinates "{{\vec e}_\varphi } = (\cos \varphi ,\sin \varphi )". Northeas is the direction with "{\varphi _1} = \frac{\pi }{4}" in our system, for Southeast we've got "{\varphi _2} = - \frac{\pi }{4}" . Then we can find the displacement vector as

"\vec d = {l_1}{{\vec e}_{{\varphi _1}}} + {l_2}{{\vec e}_{{\varphi _2}}}"

where "{l_1}" is the travelled distance along northeast direction and "{l_2}" is the travelled distance along the southeast direction. The shortest distance is the absolute value of "{\vec d}" . So let's find the vector


"\vec d = 3[{\text{km}}] \cdot (\cos \frac{\pi }{4},\sin \frac{\pi }{4}) + 4[{\text{km}}](\cos ( - \frac{\pi }{4}),\sin ( - \frac{\pi }{4}))"

Perform thecalculations

"\vec d = 3[{\text{km}}] \cdot (\frac{1}{{\sqrt 2 }},\frac{1}{{\sqrt 2 }}) + 4[{\text{km}}](\frac{1}{{\sqrt 2 }}, - \frac{1}{{\sqrt 2 }})"

"\vec d = \frac{1}{{\sqrt 2 }}(3,3)[{\text{km}}] + \frac{1}{{\sqrt 2 }}(4, - 4)[{\text{km}}] = \frac{1}{{\sqrt 2 }}(7, - 1)[{\text{km}}]"

And find the absolute value


"\left| {\vec d} \right| = \frac{1}{{\sqrt 2 }}\sqrt {{7^2} + {{( - 1)}^2}} [{\text{km}}] = \frac{1}{{\sqrt 2 }}\sqrt {{7^2} + {{( - 1)}^2}} [{\text{km}}] = \frac{{5\sqrt 2 }}{{\sqrt 2 }}[{\text{km}}] = 5[{\text{km}}]"

So the "\left| {\vec d} \right| = 5[{\text{km}}]" is the answer.